A sample survey on the average total yearly expenditure included 150 students of a certain university. The mean total expenditure per student per year for the sample was 3,000 with a standard deviation of 500. How likely is it that the students spend an average of 3500 per year as claimed by a parent at .01 significant level. 


1
Expert's answer
2021-09-28T01:22:57-0400

"n=150 \\

\bar{x}=3000 \\

s=500 \\

H_0: \mu = 3500 \\

H_1: \mu ≠3500 \\

α=0.01"

When dealing with a large sample of size n >30 from a population that need not be normal but has a finite variance, we can use the central limit theorem to justify using the test for normal populations. Even when "\sigma^2" is unknown we can approximate its value with "s^2" in the computation of the test statistic.

Test-statistic:

"Z= \frac{\bar{x} - \mu}{s/ \sqrt{n}}"

"\bar{x}" = sample mean

"\mu" = population mean

s = sample standard deviation

n = sample size

"Z = \frac{3000-3500}{500 / \sqrt{150}} = \frac{-500}{40.85} = -12.24"

Two-tailed test. Reject H0 if Z≤ -2.575 or Z ≥ 2.575.

"Z = -12.24 < Z_{tab} = -2.575"

We reject H0.

There is NOT enough evidence, that the students spend an average of 3500 per year as claimed by a parent at 0.01 significant level.